# Social welfare optimisation in well-mixed and structured populations

### v1

@@ 4.2 Numerical Simulation for well–mixed populations
    As the intensity of selection approaches neutrality ($\beta\rightarrow 0^{+}$), the expected Social Welfare is observed to converge to a constant value, rendering the objective to maximize $SW$ effectively redundant as its value does not change with $\theta$. For sufficiently high values of $\beta$, the Social Welfare exhibits a unique optimum: a local maximizer that also constitutes the global maximizer for $SW$. This $SW$ maximizer is found to coincide with the global maximizer for Cooperation Frequency and the local minimiser of Institution Cost on the interval $\theta^{0},+\infty)$. The critical point $\theta^{0}$ is defined theoretically as $\theta^{0}=\log u_{0}/\beta-\delta$, where $u_{0}$ is the unique positive root of the polynomial $P(u)$.
@@ 4.3 Agent–based simulations for structured populations
As described in Section 3.3.1, we run simulations for a population of size $10000$ in the lattice graph setup, so at beginning of each round, each agent plays with at most 4 other neighbouring agents. After each agent gets their payoff result, the external institution steps in and invest each cooperator an amount $\theta$ corresponding to either the POP or the NEB scheme, thereby influencing the decision of the agents to choose which behaviour to do in the next game round. To ensure comparability, all following simulations assume the normalized payoff matrix for Prisoner’s Dilemma game and all other settings the same as [han2018fostering]. A notable addition in the following results compared to the seminal work is the efficiency coefficient $a$ that denotes how effective the incentivisation is. If not specified in the experiments then $a=1$ which means that the social welfare is equal to the population payoff.
@@
Since $a=1$ means that the social welfare is equal to the population payoff, varying the efficiency in relation with the different $p_{C}$ and $n_{C}$, combined with the different $\theta$ could yield more insightful information. Figure 4 presents the POP intervention method across different $a$, $\theta$ and $p_{C}$. For the sake of presentation, we only plot in cases where $p_{C}/Z\geq 0.9$ which ensures that the population converges around the state where at least 90% of the population is cooperative. At $p_{C}/Z=1$, the result is predictable where with $a\leq 1$, as the cost decreases, the social welfare increases, where it is the other way around when $a>1$ (an investment into the population yields better return than the actual spending, or another way to put it is the perceived cost is higher than the actual cost), in which case as the cost increases, the social welfare also increases. So for POP, this indicates where minimising cost does not result in maximising social welfare.
@@
For NEB intervention strategy, we compare the optimal investment cost $\theta$ for minimising the total cost and boosting the cooperation frequency in population and the one for maximising social welfare in order to prove that they are not the same. From figure 5, we can observe how NEB obtain the objective in three cases of the cost effective coefficient: $a=0.5,a=1$, and $a=1.5$ on the varying value of $\theta$. We can see that the pattern for NEB to successfully reach the highest frequency of cooperation is similar to what had been explored in [han2018fostering]. The more additional important observation from the first column (the social welfare) is that the social welfare when $n_{C}=4$ is optimal when the cost is not efficient, this remain the same in all three values of $a$. Moreover, with $n_{C}=3$, the lower value of $a$, the more value of social welfare added (the orange colour changes from light to dark in the 3/K column when $2.0<\theta<3.5$ of the first diagram). At those of $\theta$ is less cost–efficient.
For clearer observation, as shown in figure 6, with $n_{C}=3$, the optimal $\theta$ for cost–efficiency is approximately 2.6 while that for social welfare optimisation is 3.0, the similar pattern can also be seen when $n_{C}=4$ (these $\theta$ value are determined when the number of cooperators in the population is greater than 90% of the total). The gap between two optimal $\theta$ state that when we try to find the efficient value of $\theta$ for concurrently minimising the cost and boosting the frequency of cooperation, the maximal value of social welfare is not obtained. In addition, at the $\theta$ to reach the maximal value of social welfare, the cost at that point is always higher than the optimal cost in both context of $n_{C}$. This proves the hypothesis that optimizing both cooperation frequency and cost does not take the social welfare to the efficient state.
@@ 5 Conclusions and Future Work
In this work, we have approach the optimisation of social welfare in well–mixed and structured populations under an external institutional investment from two angles, one using mathematical and numerical analysis and the other using agent–based simulation, to address the research question of whether optimising intervention cost means optimising social welfare. Both angles reach a consensus where the $\theta$ for optimising the intervention cost is different from optimising the social welfare. From the mathematical perspective, we have surveyed the objective function with incentive cost reaching infinity. We have found an approximation algorithm for the optimal solution in the inefficient transfer case of the reward mechanism and used numerical method to simulate some examples on the correlation between the bi–objective for minimising cost and increasing cooperation and maximising social welfare. We also conducted agent–based simulation on square lattice structured network of population to observe and analyse the difference in the $\theta$ value to take the desired configuration in both global in local interference strategies. The empirical results show that there exists a gap between them, thus proving our hypothesis about the capability of multi objectives on the optimal social welfare.
For future work, we will explore the complete behaviour of the social welfare objective function in $\mathbb{R}^{+}$ and try to find an algorithm to approximate the solution for the case of zero–sum transfer of the reward mechanism. In addition, we can try to run simulation on multiple players game like Public Goods Game to study the pattern of different criteria (total cost, social welfare) and take a deeper observation into the affect of network structure on the objectives.

### v2

@@ 4.2 Numerical Simulation for well–mixed populations
    As the intensity of selection approaches neutrality ($\beta\rightarrow 0^{+}$), the expected Social Welfare is observed to converge to a constant value, rendering the objective to maximize $SW$ effectively redundant as $SW$’s value does not change with $\theta$. For sufficiently high values of $\beta$, the Social Welfare exhibits a unique optimum: a local maximizer that also constitutes the global maximizer for $SW$. This $SW$ maximizer is found to coincide with the global maximizer for Cooperation Frequency and the local minimizer of Institution Cost on the interval $\theta^{0},+\infty)$. The critical point $\theta^{0}$ is defined theoretically as $\theta^{0}=\log u_{0}/\beta-\delta$, where $u_{0}$ is the unique positive root of the polynomial $P(u)$.
@@ 4.3 Agent–based simulations for structured populations
As described in Section 3.3.1, we run simulations for a population of size $10000$ in the lattice graph setup, so at beginning of each round, each agent plays with at most 4 other neighbouring agents. After each agent gets their payoff result, the external institution steps in and invest each cooperator an amount $\theta$ corresponding to either the POP or the NEB scheme, thereby influencing the decision of the agents to choose which behaviour to do in the next game round. To ensure comparability, all following simulations assume the normalized payoff matrix for Prisoner’s Dilemma game and all other settings the same as [han2018fostering]. A notable addition in the following results compared to the previous work is the efficiency coefficient $a$ that denotes how effective the incentivisation is. Unless otherwise specified in the experiments, we set $a=1$, so the social welfare equals the population payoff.
@@
Since $a=1$ means that the social welfare is equal to the population payoff, varying the efficiency in relation with the different $p_{C}$ and $n_{C}$, combined with the different $\theta$ could yield more insightful information. Figure 4 presents the POP intervention method across different $a$, $\theta$ and $p_{C}$. For the sake of presentation, we only plot in cases where $p_{C}/Z\geq 0.9$ which ensures that the population converges around the state where at least 90% of the population is cooperative. At $p_{C}/Z=1$, the result is predictable where with $a\leq 1$, as the cost decreases, the social welfare increases, where it is the other way around when $a>1$ (an investment into the population yields better return than the actual spending, or another way to put it is the perceived cost is higher than the actual cost), in which case as the cost increases, the social welfare also increases. So for POP, this indicates where minimizing cost does not result in maximising social welfare.
@@
For NEB intervention strategy, we compare the optimal investment cost $\theta$ for minimizing the total cost and boosting the cooperation frequency in population and the one for maximising social welfare in order to prove that they are not the same. From figure 5, we can observe how NEB obtain the objective in three cases of the cost effective coefficient: $a=0.5,a=1$, and $a=1.5$ on the varying value of $\theta$. We can see that the pattern for NEB to successfully reach the highest frequency of cooperation is similar to what had been explored in [han2018fostering]. The more additional important observation from the first column (the social welfare) is that the social welfare when $n_{C}=4$ is optimal when the cost is not efficient, this remain the same in all three values of $a$. Moreover, with $n_{C}=3$, the lower value of $a$, the more value of social welfare added (the orange colour changes from light to dark in the 3/K column when $2.0<\theta<3.5$ of the first diagram). At those of $\theta$ values is less cost–efficient.
For clearer observation, as shown in figure 6, with $n_{C}=3$, the optimal $\theta$ for cost–efficiency is approximately 2.6 while that for social welfare optimisation is 3.0, the similar pattern can also be seen when $n_{C}=4$ (these $\theta$ value are determined when the number of cooperators in the population is greater than 90% of the total). The gap between two optimal $\theta$ state that when we try to find the efficient value of $\theta$ for concurrently minimizing the cost and boosting the frequency of cooperation, the maximal value of social welfare is not obtained. In addition, at the $\theta$ to reach the maximal value of social welfare, the cost at that point is always higher than the optimal cost in both context of $n_{C}$. This proves the hypothesis that optimizing both cooperation frequency and cost does not take the social welfare to the efficient state.
@@ 5 Conclusions and Future Work
In this work, we have approached the optimisation of social welfare in well-mixed and structured populations under external institutional investment from two complementary perspectives: one using mathematical and numerical analysis, and the other employing agent-based simulation. Our goal is to investigate whether optimising intervention costs leads to an optimisation of social welfare.
Both perspectives reach a consensus that the $\theta$ for optimising intervention costs differs from that for optimising social welfare. From the mathematical analysis in well-mixed populations, we have examined the objective function with incentive costs approaching infinity. We have developed an approximation algorithm for the optimal solution in the case of inefficient transfer within the reward mechanism and employed numerical methods to simulate examples illustrating the correlation between the bi-objectives of minimising costs, increasing cooperation, and maximising social welfare. Moreover, we conducted agent-based simulations on a square lattice structured network of the population to observe and analyse the differences in the $\theta$ values needed to achieve the desired configurations under both global and local interference strategies. The simulation results indicate a gap between the two, thereby supporting our hypothesis regarding the effectiveness of multi-objective approaches on optimal social welfare.
As future work, we will investigate the complete behaviour of the social welfare objective function in $\mathbb{R}^{+}$ and seek to develop an algorithm that approximates the solution for the case of zero-sum transfers within the reward mechanism. Additionally, we plan to conduct simulations on multiplayer games, such as the Public Goods Game, to examine the patterns of various criteria (total cost, social welfare) and gain deeper insights into the effects of network structure on these objectives.